Capital University, a mid-sized university located in a capital city has approximately 10,000 students studying in four departments. While the University’s goal is not to make a profit, it is also important that the University does not operate at a loss. The only source of revenue is tuition and government grants. The government gives the University $14,000 a year for each full-time student.

Faculty:

Arts

Science

Engineering

Business

Number of Courses

120

200

150

150

Number of Students

4,800

2,000

600

2,700

Other Information:

  • Each Course is only offered once each year
  • Each student is enrolled in 10 classes per year
  • Each faculty member is paid $120,000 per year and teaches 6 courses per year
  • Total student support costs are $139,992,700 per year. These costs include student services and administration. Each student generally requires an equal amount of support.
  • Building costs are $12,400,000 per year. Costs are equal for each class.
  • Other costs $17,000,000 (Science labs $10million, Business trading room $2million, Engineering labs $5million).

REQUIRED

  1. The Dean has asked you to use this information to determine how much tuition should be charged assuming the same amount of tuition should be charged to each student and the University breaks even.
  2. Using Activity Based Costing, calculate how much tuition should be charged to students in each faculty.
Answer #1

Total courses= 120 + 200 +150 + 150 = 620

Total teachers required = 620/6 =103 (rounded off)

Total cost of teacher = 103 x120000 =12360000

Total stydent support cost =139992700

Building cost =12400000

Other cost =17000000

Total cost =12360000 +139992700 +12400000 +17000000 =181752700

Grants received =10100 x 14000 = 141400000

Cost to be charged to students = 181752700 - 141400000 =40352700

Tution fee to be charged per student = 41752700/10100 = $3995.32